# Mind the Gap: Removing the Discretization Gap in Differentiable Logic Gate Networks

### v1

@@ Appendix B Gumbel Smoothing
The proof of Lemma 1 depends on the translation invariance of the softmax.
@@
Restating lemma lemma 1, we write:
Lemma 3 (Gumbel-Smoothing).
@@ Appendix C Distribution over Logic Gates
We look in Figure 10 at the distribution of the logic gates in the final network. Our first observation is that the distributions are far more uniform for Gumbel LGNs in layers 1 to 11 than for Differentiable LGNs. At the same time, we see a sharp transition for Differentiable LGNs after layer 8. This could match the results in Figure 7, indicating that neurons in Differentiable LGNs struggle to converge in all but the final layers.
@@
Figure 10: Gate distribution for the gates split by layer for a Differentiable LGN and a Gumbel LGN. Interestingly, the distribution is far more uniform for the Gumbel LGN in layers 1 to 11 than for the Differentiable LGN. In addition, the Gumbel LGN primarily has “1” gates in the final layer, which can be seen as a constant bias towards certain classes. We analyze this further in Appendix C.
@@
We see in Figure 11 the gate distribution for each of the 10 classes in the last layer. Interestingly, the distributions between Softmax and Gumbel are quite different, but the classes are nearly identical. Since almost all classes have several “1” gates, pruning these would be possible as softmax is translation invariant.
@@ Appendix D Extended Sharpness-Aware Minimization
where $p\in[1,\infty[$ is the $p$-norm used (usually $p=2$) and $\rho>0$ is a hyperparameter [3]. This optimization objective arises from the PAC-Bayesian generalization bound shown in Equation 8 [62], and Foret et al. [3] use it to show Equation 9. Below, $n=|\mathcal{S}|$, $k$ is the number of model parameters, $\mathscr{P}$ and $\mathscr{Q}$ are the prior and posterior distributions of the model parameters, respectively. Lastly, the equations hold with probability $1-\delta$.
@@ Appendix F Estimating Hessian of Loss
Computing the full Hessian of the loss function is not computationally feasible due to the quadratic scaling with the number of parameters. Instead, we use scalable stochastic methods to estimate the trace of the Hessian, which provides useful curvature information. Specifically, we focus on the trace as our Lemma 1 directly minimizes the trace of the Hessian.
@@ Appendix G Loss Surface Visualization
where $d_{1}$ and $d_{1}$ are orthogonal directions with $\alpha,\beta\in\mathbb{R}$. We choose $\alpha,\beta$ such that we probe the model in a unit circle, i.e. $(-1,1)$.
@@ H.1 Hyperparameters
we use the same hyperparameters for this experiment for Differentiable LGNs and Gumbel LGNs. Specifically, we fix the width of the network to $256K$ neurons and train the network over the depths $\{6,8,10,12\}$. We optimize the network using Adam with a learning rate of $0.01$. Furthermore, the batch size is set to $128$ and the final parameter in the GroupSum is set to $1/0.01$. This mirrors the original experimental setup of Petersen et al. [1] for CIFAR-10. Finally, we fix the $\tau=1$ parameter for the Gumbel noise in the Gumbel LGNs.
@@ Appendix I Softmax DiffLogic Performance on MNIST-like Baselines
For all datasets, we use the same model and experiment configs. [^3] Thus, the CIFAR numbers are not representative of the optimal performance. The models have six layers and a width of 64k. See Table 3 for the results.
@@
Table 3: The table shows the performance of Differentiable LGNs on many datasets. The key takeaway is that the discretization gap for MNIST-like datasets is minimal. The discretization gap is the difference between the discrete and soft performance. The numbers are averaged over five runs.
@@
In Table 4 the number of iterations per hour Softmax and Gumbel LGNs completed while training models for the results in Figure 3. We also calculate the relative difference between the two. Gumbel is slightly slower due to the noise sampling; however, as it converges much faster in terms of iterations, the net effect is that it converges $4.5$ times faster in wall-clock time.
Table 4: Iterations per hour and the relative change from Differentiable LGNs to Gumbel LGNs.

### v2

@@ Appendix B Gumbel Smoothing
The proof of Lemma˜1 depends on the translation invariance of the softmax.
@@
Restating Lemma˜1
Lemma 3 (Gumbel‐Smoothing).
@@ Appendix C Distribution over Logic Gates
We look in Figure˜12 at the distribution of the logic gates in the final network. Our first observation is that the distributions are far more uniform for Gumbel LGNs in layers 1 to 11 than for DLGNs. At the same time, we see a sharp transition for DLGNs after layer 8. This could match the results in Figure˜9, indicating that neurons in DLGNs struggle to converge in all but the final layers.
@@
Figure 12: Gate distribution for the gates split by layer for a Differentiable LGN and a Gumbel LGN. Interestingly, the distribution is far more uniform for the Gumbel LGN in layers 1 to 11 than for the Differentiable LGN. In addition, the Gumbel LGN primarily has “1” gates in the final layer, which can be seen as a constant bias towards certain classes. We analyze this further in Appendix˜C.
@@
We see in Figure˜13 the gate distribution for each of the 10 classes in the last layer. Interestingly, the distributions between Softmax and Gumbel are quite different, but the classes are nearly identical. Since almost all classes have several “1” gates, pruning these would be possible as softmax is translation invariant.
@@ Appendix D Extended Sharpness-Aware Minimization
where $p\in[1,\infty[$ is the $p$-norm used (usually $p=2$) and $\rho>0$ is a hyperparameter [15]. This optimization objective arises from the PAC-Bayesian generalization bound shown in Equation˜8 [73], and Foret et al. [15] use it to show Equation˜9. Below, $n=|\mathcal{S}|$, $k$ is the number of model parameters, $\mathscr{P}$ and $\mathscr{Q}$ are the prior and posterior distributions of the model parameters, respectively. Lastly, the equations hold with probability $1-\delta$.
@@ Appendix F Estimating Hessian of Loss
Computing the full Hessian of the loss function is not computationally feasible due to the quadratic scaling with the number of parameters. Instead, we use scalable stochastic methods to estimate the trace of the Hessian, which provides useful curvature information. Specifically, we focus on the trace as our Lemma˜1 directly minimizes the trace of the Hessian.
@@ Appendix G Loss Surface Visualization
where $d_{1}$ and $d_{1}$ are orthogonal directions with $\alpha,\beta\in\mathbb{R}$. We choose $\alpha,\beta$ such that we probe the model in a unit circle, i.e. (1,1)(-1,1).
@@ H.1 Hyperparameters
we use the same hyperparameters for this experiment for DLGNs and Gumbel LGNs. Specifically, we fix the width of the network to $256K$ neurons and train the network over the depths $\{6,8,10,12\}$. We optimize the network using Adam with a learning rate of $0.01$. Furthermore, the batch size is set to $128$ and the final parameter in the GroupSum is set to $1/0.01$. This mirrors the original experimental setup of Petersen et al. [7] for CIFAR-10. Finally, we fix the $\tau=1$ parameter for the Gumbel noise in the Gumbel LGNs.
@@ Appendix I Softmax DiffLogic Performance on MNIST-like Baselines
For all datasets, we use the same model and experiment configs. [^2] Thus, the CIFAR numbers are not representative of the optimal performance. The models have six layers and a width of 64k. See Table˜3 for the results.
@@
Table 3: The table shows the performance of DLGNs on many datasets. The key takeaway is that the discretization gap for MNIST-like datasets is minimal. The discretization gap is the difference between the discrete and soft performance. The numbers are averaged over five runs.
@@
In Table˜4 the number of iterations per hour Softmax and Gumbel LGNs completed while training models for the results in Figure˜3. We also calculate the relative difference between the two. Gumbel is slightly slower due to the noise sampling; however, as it converges much faster in terms of iterations, the net effect is that it converges $4.5$ times faster in wall-clock time.
Table 4: Iterations per hour and the relative change from DLGNs to Gumbel LGNs.